OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-8,-10_ and is perpendicular to -8-3y=-16

OpenStudy (aroub):

do u mean -8x-3y=-16 or -8-3y=-16x ???

OpenStudy (anonymous):

the first one

OpenStudy (aroub):

okay! :)

OpenStudy (aroub):

y=mx+b ,m=3/8 , (-8,-10) -10=3/8(-8)+b -10=-3+b -7=b so the equation is y=3/8x-7

OpenStudy (aroub):

this question is from a book or online test or ... ???

OpenStudy (anonymous):

online test